OpenStudy (anonymous):

Which equation produces a line that is perpendicular to the line represented by the function below Y=2/5x+9 A)5x+2y=4 B) 2x-5y=8 C) 5x-2y=-3 D) 2x+5y=-7

OpenStudy (snuggielad):

So, this one wants to be tricky. They give you slope intercept and you have to find it in standard. I will start out answering by asking you, Do you know the difference between slope intercept and standard?

OpenStudy (snuggielad):

@ItsStephanie ?

OpenStudy (anonymous):

Yes

OpenStudy (snuggielad):

How do you convert a standard problem to slope intercept?

OpenStudy (anonymous):

Well I know what they both look like but not how to convert

OpenStudy (snuggielad):

So if I tell you the slope of the perpendicular is going to be \(-\frac{ 5 }{ 2 }\) what would be your choice?

OpenStudy (anonymous):

Between the two?

OpenStudy (snuggielad):

the first one

OpenStudy (anonymous):

Slope-intercept form

OpenStudy (anonymous):

@SnuggieLad im completely lost actually

OpenStudy (loser66):

The given equation is y = (2/5) x +9 Hence the perpendicular line will have the slope is \(-\dfrac{5}{2}\) Then, the required line will have the equation : \(y = -\dfrac{5}{2}x +b\) \(2y = -5x+2b\\2y+5x =2b\) That means the right hand side must be an even number. You have 2 options whose the left hand side are the same \(2y+5x\) But only one whose the right side is an even number. That is ....... ????